OpenStudy (anonymous):

I am to find the series for y''+2xy=0

OpenStudy (abb0t):

use the guess \[y = \sum_{n=0}^{\infty} C_nx^n\]

OpenStudy (abb0t):

find the first and second derivative for \(y\) which is \[y'' = \sum_{n=2}^{\infty}n(n-1)C_nx^{n-2}\]

OpenStudy (abb0t):

substitute into your given differential equation to get: \[\sum_{n=2}^{\infty}n(n-1)C_nx^{n-2}+2x \sum_{n=0}^{\infty}C_nx^n = 0\] distribute the \(2x\) inside the sum \[\sum_{n=2}^{\infty}n(n-1)C_nx^{n-2}+2\sum_{n=0}^{\infty}C_nx^{n+1}=0\]

OpenStudy (abb0t):

Are you with me so far?

OpenStudy (anonymous):